# Longhorn: State Space Models are Amortized Online Learners

### v3

@@ 1 Introduction
In this work we propose one potential principle. We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling. From this perspective:
@@
Figure 2: (left) Most existing sequence models consist of channel and sequence mixing layers. The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $s_{t}$, which is then passed to later layers for sequence modeling. (middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows a certain dynamics that optimizes an online learning objective. The recurrent update of $s$ can then be derived by solving the online objective in closed-form or by conducting a gradient step. (right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$. The closed-form solution to the online regression problem corresponds to the implicit online learning update, where $\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$. Please see the details in Section 3.
As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems. The hope is that by selecting the right objective, one can save parameters / computation without sacrificing performance. In addition, the proposed online learning perspective might lead to better understanding of what SSM layers are doing in a large model. Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective. We provide the objectives for some existing SSMs in Table 1.
@@ 2 Background
Deep SSMs replace the SA layers with SSM layers. Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023). But in both cases, the sequence mixing is done by the SSM module and the channel mixing is done by the channel-wise MLP. Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of a SSM block for sequence mixing (red), and a MLP block for channel mixing (blue).
@@
Existing variants of SSMs mainly differs in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence. A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed be $m$ times of the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence. This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging. A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large. This makes it necessary to impose certain low dimensional structure in $A_{t},B_{t}$, which is what differentiate the existing designs of SSM. In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
@@
where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product.
In practice, one can choose to use $h$ heads as in the multi-head attention to save some computation, where both the $m$ and $d$ dimension are divided into $h$ groups and each group performs its own LA variant. The outer product becomes $\mathcal{O}(h*m/h*d/h=md/h)$. But then the effective size of $S_{t}$ also shrinks to $md/h$.
@@
We can see that in Mambda both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
@@
In Mamba and the linear attention variants, the outer product serves a critical rule of lifting vectors to matrices. The recent Griffin architecture abandons the outer product and performs pure elementwise product:
@@
where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$. This yields smaller memory states, but in practice Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
@@
where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting. In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with outer product, which makes it more similar to a LA variant.
@@ 3 An Online Learning Perspective For Sequence Mixing
As demonstrated in the previous section, the design of a state-space model (SSM) depends on the specific selection of $(A_{t},B_{t},*)$, a process that is both intricate and somewhat artisanal. In this section, we propose to streamline SSM design through a online learning perspective. The main idea is to treat the SSM layers as learning modules that learn to compress information along the sequence dimension. From this perspective, the SSM layers are learning to learn, such that at the inference time, these layers are still learning (compressing) new information online.
@@ 3.1 SSM as Online Learning
We advocate to view the recurrence of SSM as solving an online learning problem. In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$. The goal is to minimize
@@
Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013). Formally, an online convex programming algorithm update $s_{t}$ by solving a regularized cost function:
@@
where $\beta_{t}\in\mathbb{R}^{+}$ and $D_{\phi}$ is a discrepancy measure, often a Bregman divergence induced by the convex function $\phi$ (e.g., when $\phi(x)=\frac{1}{2}\left\lVert x\right\rVert^{2}$, $D_{\phi}(s,s_{t-1})=\frac{1}{2}||s-s_{t-1}||^{2}$). Here the first term ensures the updated $s$ will be close to the previous $s_{t-1}$, so the agent suffers less from catastrophic forgetting, while the second term ensures the agent is incorporating new knowledge from minimizing the new loss $\ell_{t}(s)$. Hence, $\beta_{t}$ controls the trade-off between stability and plasiticity.
@@ 3.2 SSM as Online Learning and The Longhorn Architecture
Under the online learning framework, the design of SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 5. This provides a unified framework for the existing SSM variants. We summarize in Table 1 the online learning interpolation of a number of existing SSM architectures.
In this work, we explore a highly simplified and natural design called Longhorn guided the online principle (see the last row of Table 1). In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is higher dimensional). Then, we propose to use the following OCP objective:
$s_{t}=\operatorname*{arg\,min}_{s}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
Here, $k_{t}(x_{t})\in\mathbb{R}^{d}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer. $\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate. For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity. For simplicity, we will refer to them as $k_{t},\beta_{t}$ from now on. From a high-level, Equation 6 is solving an online prediction problem, also known as the associative memory problem (a supervised learning problem), where the goal is to memorize all $(k_{t},x_{t})$ pairs (learn a mapping from $k_{t}$ to $x_{t}$) such that given a key (input) $k$ the model can retrieve (predict) its corresponding value (label) $x$.
Fortunately, this simple objective gives a closed form solution for $s_{t}$ (a.k.a., the implicit online learning method (Kulis & Bartlett, 2010)).
@@
As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting parallel scan. As a result, in practice, we use the diagonal approximation $1-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$. The final SSM recurrence therefore becomes:
@@
When $x_{t}\in\mathbb{R}^{d}$ is a higher dimensional vector, following the Mamba design, we use $d$ parallel SSMs. As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
@@
The final architecture of Longhorn follows Mamba strictly (Figure 3), and we only replace the SSM block with the Longhorn’s recurrence. We also provide an efficient CUDA kernel for it. The full inference time algorithm is provided in Algorithm 1. One can compare Equation 9 to Equation 3 and other SSM in Section 2. Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forget gate is naturally derived from the key vector, i.e., $(1-k^{\odot 2}\otimes\Delta_{t})$.
Advantages of Longhorn: 1. While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem. 2. Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba). The forgetting is naturally linked to the key vector $k_{t}$ through the derivation. This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$. However, Longhorn demonstrates better performance even with fewer parameteres than Mamba (See Figure 1 (left), Table 3, Table 4). 3. The closed-form solution in Equation 7 does not need any specific initialization. In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$. Comparing to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA). Therefore, it is as fast as existing SSMs. 4. Unlike DeltaNet, which, as noted in Yang et al. (2024), cannot extrapolate beyond the training context, Longhorn demonstrates the capability to successfully extrapolate to contexts up to 16x longer at inference time (See Figure 1 (right)).
@@ 6 Conclusion and Future Work
In this work, we introduce a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective. We propose a straightforward online regression objective, adopting its implicit closed-form update to define our model, which we refer to as Longhorn. Notably, Longhorn is highly parallelizable and achieves state-of-the-art performance in both synthetic sequence modeling and language modeling tasks. For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware. Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improve performance. We anticipate similar benefits for Longhorn.

### v4

@@ 1 Introduction
In this work, we propose one potential principle. We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling. From this perspective:
@@
Figure 2: (left) Most existing sequence models consist of channel and sequence mixing layers. The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $s_{t}$, which is then passed to later layers for sequence modeling. (middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows certain dynamics that optimize an online learning objective. The recurrent update of $s$ can then be derived by solving the online objective in closed form or by conducting a gradient step. (right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$. The closed-form solution to the online regression problem corresponds to the implicit online learning update, where $\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$. Please see the details in Section 3.
As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems. The hope is that by selecting the right objective, one can save parameters/computation without sacrificing performance. In addition, the proposed online learning perspective might lead to a better understanding of what SSM layers are doing in a large model. Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective. We provide the objectives for some existing SSMs in Table 1.
@@ 2 Background
Deep SSMs replace the SA layers with SSM layers. Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023). But in both cases, the sequence mixing is done by the SSM module, and the channel mixing is done by the channel-wise MLP. Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
@@
Existing variants of SSMs mainly differs in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence. A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed be $m$ times of the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence. This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging. A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large. This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is what differentiates the existing designs of SSM. In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
@@
where $S_{t}\in\mathbb{R}^{d\times m}$, $v(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ are linear mappings of $x_{t}$, and $\otimes$ denote the outer product. In practice, one can use $h$ heads as in the multi-head attention to save some computation, where the $m$ and $d$ dimensions are divided into $h$ groups and each group performs its own LA variant. The outer product complexity reduces to $\mathcal{O}(h*m/h*d/h=md/h)$. But then the effective size of $S_{t}$ also shrinks to $md/h$.
@@
In Mamba, both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
@@
In Mamba and the linear attention variants, the outer product serves as a critical rule of lifting vectors to matrices. The recent Griffin architecture abandons the outer product and performs pure elementwise product:
@@
where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$. This yields smaller memory states, but in practice, Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
@@
where all the vectors are of size $\mathbb{R}^{d}$, and $w>0$ is a trainable weight for controlling the forgetting. In the most recent RWKV version (Peng et al., 2024), the denominator $z_{t}$ is removed, and the elementwise product is replaced with the outer product, which makes it more similar to an LA variant.
@@ 3 An Online Learning Perspective For Sequence Mixing
As demonstrated in the previous section, designing a state-space model (SSM) depends on the specific selection of $(A_{t},B_{t},*)$, which is intricate and somewhat artisanal. In this section, we propose to streamline SSM design through an online learning perspective. The main idea is to treat the SSM layers as learning modules that learn to compress information along the sequence dimension. From this perspective, the SSM layers are learning to learn, such that during the inference time, these layers are still learning (compressing) new information online.
@@ 3.1 SSM as Online Learning
We advocate viewing the recurrence of SSM as solving an online learning problem. In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$. The goal is to minimize
@@
Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013). Formally, an online convex programming algorithm updates $s_{t}$ by solving a regularized cost function:
@@
where $\beta_{t}\in\mathbb{R}^{+}$ and $D_{\phi}$ is a discrepancy measure, often a Bregman divergence induced by the convex function $\phi$ (e.g., when $\phi(x)=\frac{1}{2}\left\lVert x\right\rVert^{2}$, $D_{\phi}(s,s_{t-1})=\frac{1}{2}||s-s_{t-1}||^{2}$). Here the first term ensures the updated $s$ will be close to the previous $s_{t-1}$, so the agent suffers less from catastrophic forgetting, while the second term ensures the agent is incorporating new knowledge from minimizing the new loss $\ell_{t}(s)$. Hence, $\beta_{t}$ controls the trade-off between stability and plasticity.
@@ 3.2 SSM as Online Learning and The Longhorn Architecture
Under the online learning framework, the design of SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 5. This provides a unified framework for the existing SSM variants. We summarize in Table 1 the online learning interpretation of several existing SSM architectures.
In this work, we explore a highly simplified and natural design called Longhorn guided by the online principle (see the last row of Table 1). In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is a vector). We want to recurrently update hidden states $\{s_{t}\}_{t}$, where $s_{t}\in\mathbb{R}^{m}$ is a vector that aims to summarize the information up to time $t$. We posit the following OCP objective for updating $s_{t}$:
$s_{t}=\operatorname*{arg\,min}_{s\in\mathbb{R}^{m}}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
Here, $k_{t}(x_{t})\in\mathbb{R}^{m}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer. $\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate. For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity. For simplicity, we will write $k_{t}=k_{t}(x_{t})$ and $\beta_{t}=\beta_{t}(x_{t})$ from now on.
From a high-level, Equation 6 is solving an online prediction problem of learning a weight $s$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx s^{\top}k_{t}$. It is a supervised formulation of the associative memory problem of memorize $(k_{t},x_{t})$ pairs by learning a mapping from $k_{t}$ to $x_{t}$, such that given a key (input) $k_{t}$ the model can retrieve (predict) its corresponding value (label) $x_{t}$.
Fortunately, this simple objective gives a closed-form solution for $s_{t}$, which coincides with the implicit online learning method (Kulis & Bartlett, 2010).
@@
As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting a parallel scan. As a result, in practice, we use the diagonal approximation $1_{m}-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$ and $1_{m}$ is the $m$-dimensional vector of all ones. The final SSM recurrence therefore becomes:
@@
Following the Mamba design, when $x_{t}\in\mathbb{R}^{d}$ is a vector, we use $d$ parallel SSMs. As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
@@
The final architecture of Longhorn follows Mamba strictly (Figure 3), and we only replace the SSM block with the Longhorn’s recurrence. We also provide an efficient CUDA kernel for it. The full inference-time algorithm is provided in Algorithm 1. One can compare Equation 9 to Equation 3 and other SSM in Section 2. Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forget gate is naturally derived from the key vector, i.e., $(1-k^{\odot 2}\otimes\Delta_{t})$.
Advantages of Longhorn: 1. While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem. 2. Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba). The forgetting is naturally linked to the key vector $k_{t}$ through the derivation. This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$. However, Longhorn demonstrates better performance even with fewer parameters than Mamba (See Figure 1 (left), Table 3, Table 4). 3. The closed-form solution in Equation 7 does not need any specific initialization. In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$. Compared to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA). Therefore, it is as fast as existing SSMs. 4. Unlike DeltaNet, which, as noted in Yang et al. (2024), cannot extrapolate beyond the training context, Longhorn demonstrates the capability to successfully extrapolate to contexts up to 16x longer at inference time (See Figure 1 (right)).
@@ 6 Conclusion and Future Work
In this work, we introduce a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective. We propose a straightforward online regression objective, adopting its implicit closed-form update to define our model, which we refer to as Longhorn. Notably, Longhorn is highly parallelizable and achieves state-of-the-art performance in both synthetic sequence modeling and language modeling tasks. For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware. Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance. We anticipate similar benefits for Longhorn.